Show that sinx and cosx are a bound variation on a finite interval


1
Expert's answer
2021-05-24T17:19:34-0400

"\text{We know , if a function is Lipschitz function then it is bounded variation.}\newline 
\text{Given,} f_1(x)=sinx. \,\\
|f_1(x_1)-f_1(x_2)|\newline
=|sin(x_1)-sin(x_2)|\hspace{1cm}(\text{By trigonometry identity}\newline
=|2 cos(\frac{x_1+x_2}{2}) sin(\frac{x_1-x_2}{2})|\newline
\leq|2|.1.|\frac{x_1-x_2}{2}|\newline
=|x_1-x_2|\newline \text{
we get k=1.}\newline \text{
Therefore, the given function is lipschitz function.}\newline 
\text{This implies, the given function is bounded variation.}
\text{Given,} f_2(x)=cosx.\\
|f_1(x_1)-f_1(x_2)|\newline
=|cos(x_1)-cos(x_2)|\hspace{1cm}(\text{By trigonometry identity}\newline
=|-2 sin(\frac{x_1+x_2}{2}) sin(\frac{x_1-x_2}{2})|\newline
\leq|2|.1.|\frac{x_1-x_2}{2}|\newline
=|x_1-x_2|\newline \text{
we get k=1.}\newline \text{
Therefore, the given function is lipschitz function.}\newline 
\text{This implies, the given function is bounded variation.}"